# Fluctuations in random field Ising models

### v1

@@ 3 Proof of main results
    where the last line uses Lemma 3.4 on noting that $1-\lambda_{n}\upsilon_{n}\in[1-\rho,1/(1-\rho)]$. Part (b) now follows on plugging-in the bound in part (a).
@@
$\operatorname{\mathbb{E}}(\upsilon_{n}-\upsilon)^{2}$ $=O(\|\mathbf{q}\|_{\infty}^{2}),\quad\operatorname{\mathbb{E}}R_{1n}=O(\alpha_{n}+\|{\bm{\epsilon}}\|^{2}),\quad\operatorname{\mathbb{E}}R_{2n}=O(n\alpha_{n}),$ (4.1)
$\quad\operatorname{\mathbb{E}}R_{3n}$ $=O(n\alpha_{n}^{2}),\quad\operatorname{\mathbb{E}}R_{4n}=O(\sqrt{\operatorname{\alpha_{\mathnormal{n}}}}+\|\mathbf{q}\|_{\infty}),\quad\operatorname{\mathbb{E}}|{\bm{\epsilon}}^{\top}\Psi^{\prime}({\bf c})|=O(\|\bm{\epsilon}\|).$
We defer the proof of (4.1) to the end of the proof.
@@
    Using 2.4, on the set $\upsilon_{n}\geq\upsilon/2$ the first term in the RHS of (4.2) can be bounded as follows:
    $d_{KS}(T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}),W_{n}\mid{\bf c})\lesssim\sqrt{R_{1n}}+\sqrt{\alpha_{n}R_{2n}}+\sqrt{R_{3n}}+\sqrt{n}\alpha_{n}+\|{\bf q}\|_{\infty}+\|{\bm{\epsilon}}\|_{2}=:S^{(1)}_{n}.$ (4.3)
@@
    where we use Lemma 3.4, along with the fact that $1-\lambda\upsilon_{n}\in[1-\rho,1/(1-\rho)]$. The desired conclusion follows by combining the above three displays, on using (4.1) to note that
    $\operatorname{\mathbb{E}}[S^{(1)}_{n}]\lesssim_{F}\textsc{Err}(\mathbf{A}_{n},\mathbf{q}),\quad\text{ and }\quad\operatorname{\mathbb{P}}(\upsilon_{n}\geq\upsilon/2)\to 1.$
@@
    $X=\mathbf{q}^{\top}\left(\operatorname{\mathnormal{\bm{\sigma}}}-\frac{\Psi^{\prime}(\mathbf{c})}{1-\lambda\upsilon}\right),\quad\tau=\frac{1}{1-\lambda\upsilon}\in[1-\rho,1/(1-\rho)],\quad a=1,\quad b=\mathbf{q}^{\top}\left(\mathbf{u}-\frac{\Psi^{\prime}(\mathbf{c})}{1-\lambda\upsilon_{n}}\right)$
@@
    where in the last inequality we use 2.5 part (b), and restrict ourselves to the set $\upsilon_{n}\geq\upsilon/2$, and
    $S^{(2)}_{n}:=\sqrt{R_{2n}}\left(\sqrt{R_{1n}}+\sqrt{\alpha_{n}}+\lVert\bm{\epsilon}\rVert\right)+\sqrt{R_{1n}}+\sqrt{R_{3n}}+\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}+\lVert\mathbf{q}\rVert_{\infty}+R_{4n}+\big{|}\bm{\epsilon}^{\top}\Psi^{\prime}(\mathbf{c})\big{|}.$
@@
    $\operatorname{\mathbb{E}}[S^{(2)}_{n}]\lesssim_{F}\textsc{Err}(\mathbf{A}_{n},\mathbf{q})+\sqrt{n\alpha_{n}}\|{\bm{\epsilon}}\|_{2}.$
@@
    $d_{KS}\left(\sum_{i=1}^{n}q_{i}\Big{(}\sigma_{i}-\frac{\psi^{\prime}(c_{i})}{1-\lambda\upsilon}\mid{\bf c}\Big{)},W_{\infty}\right)=O_{P}(\textsc{Err}(\mathbf{A}_{n},\mathbf{q})+\sqrt{n\alpha_{n}}\|{\bm{\epsilon}}\|_{2}),$
@@
    Proceeding to verify the second (i.e. annealed) conclusion of part (b), let $(W_{\infty},\widetilde{W}_{\infty})$ be mutually independent Gaussians, coupled in the same space as $(\operatorname{\mathnormal{\bm{\sigma}}},{\bf c})$ via an independent coupling. Then, setting $W^{*}_{n}:=\frac{1}{1-\lambda\upsilon}\sum_{i=1}^{n}q_{i}\psi^{\prime}(c_{i})$, triangle inequality gives
    $d_{KS}\left(\sum_{i=1}^{n}q_{i}\sigma_{i},W_{\infty}+\widetilde{W}_{\infty}\right)\leq$ $d_{KS}\left(\sum_{i=1}^{n}q_{i}\sigma_{i},W_{\infty}+W_{n}^{*}\right)+d_{KS}\left(W_{\infty}+W_{n}^{*},W_{\infty}+\widetilde{W}_{\infty}\right)$
    $\leq$ $\operatorname{\mathbb{E}}\left[d_{KS}\left(\sum_{i=1}^{n}q_{i}\sigma_{i},W_{\infty}+W_{n}^{*}\mid{\bf c}\right)\right]+d_{KS}\left(W_{n}^{*},\widetilde{W}_{\infty}\right),$
@@
    where the last inequality holds on the set $\upsilon_{n}\geq\upsilon/2$, and we have used ((b)). Noting that $d_{KS}$ is always bounded by $1$, the above two displays give
    $d_{KS}(\sum_{i=1}^{n}q_{i}\sigma_{i},W_{\infty}+\widetilde{W}_{\infty})\lesssim\operatorname{\mathbb{E}}S_{n}^{(2)}+\operatorname{\mathbb{E}}\left|(\upsilon_{n}-\upsilon)\sum_{i=1}^{n}q_{i}\psi^{\prime}(c_{i})\right|+\operatorname{\mathbb{P}}(\upsilon_{n}<\upsilon/2)+d_{KS}\left(W_{n}^{*},\widetilde{W}_{\infty}\right).$ (4.6)
@@
    $\operatorname{\mathbb{E}}S_{n}^{(2)}\lesssim$ $\textsc{Err}(\mathbf{A}_{n},\mathbf{q})+\sqrt{n\alpha_{n}}\|{\bm{\epsilon}}\|_{2},$
    $\operatorname{\mathbb{E}}\left|(\upsilon_{n}-\upsilon)\sum_{i=1}^{n}q_{i}\psi^{\prime}(c_{i})\right|\leq$ $\sqrt{\operatorname{\mathbb{E}}(\upsilon_{n}-\upsilon)^{2}}\sqrt{\operatorname{\mathbb{E}}\Big{[}\sum_{i=1}^{n}q_{i}\psi_{i}(c_{i})\Big{]}^{2}}\lesssim\|{\bf q}\|_{\infty}.$
@@
    $\operatorname{\mathbb{P}}(\upsilon_{n}<\upsilon/2)\leq\operatorname{\mathbb{P}}(|\upsilon_{n}-\upsilon|>\upsilon/2)\lesssim_{F}\operatorname{\mathbb{E}}|\upsilon_{n}-\upsilon|\lesssim\|{\bf q}\|_{\infty},$
    where the last bound again uses (4.1). Finally, the fourth term in the RHS of (4.6) can be bounded using the standard Berry-Esseen CLT for IID random variables, which gives
@@
Proof of (4.1). First we bound $\upsilon_{n}-\upsilon$. We use the independence of $\mathbf{c}=(c_{1},\ldots,c_{n})$ and the definition of $\upsilon=\operatorname{\mathbb{E}}[\psi^{\prime\prime}(c_{1})]>0$ to get
$\operatorname{\mathbb{E}}(\upsilon-\upsilon_{n})^{2}=\mbox{Var}(\psi^{\prime\prime}(c_{1}))\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{4}\leq\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{4}\leq\|\mathbf{q}\|_{\infty}^{2}.$ (4.7)
@@
$\operatorname{\mathbb{E}}R_{1n}$ $=\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\operatorname{\mathbb{E}}\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}(\psi^{\prime\prime}(c_{j})-\upsilon_{n})\right)^{2}$
$\leq 2\operatorname{\mathbb{E}}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}(\psi^{\prime\prime}(c_{j})-\upsilon)\right)^{2}+2\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}\right)^{2}\operatorname{\mathbb{E}}(\upsilon-\upsilon_{n})^{2}$
$=2\operatorname{Var}(\psi^{\prime\prime}(c_{1}))\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)^{2}q_{j}^{2}+2\lVert\mathbf{A}_{n}\mathbf{q}\rVert^{2}\operatorname{\mathbb{E}}(\upsilon-\upsilon_{n})^{2}$
@@
$\lVert\mathbf{A}_{n}\mathbf{q}\rVert^{2}\operatorname{\mathbb{E}}(\upsilon-\upsilon_{n})^{2}$ $\leq 2\lVert\bm{\epsilon}\rVert^{2}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{4}+2\sum_{\mathnormal{i}=1}^{\mathnormal{n}}(\lambda q_{i})^{2}q_{i}^{2}$
@@
The bounds for $R_{2n}$ and $R_{3n}$ also follow similarly by explicit computations leveraging the independence of $\mathbf{c}$ and $\operatorname{\mathbb{E}}[\psi^{\prime}(c_{1})]=0$, as follows:
$\operatorname{\mathbb{E}}R_{2n}=\sum_{i=1}^{n}\operatorname{\mathbb{E}}\left(\sum_{j=1}^{n}\mathbf{A}_{n}(i,j)\psi^{\prime}(c_{j})\right)^{2}=\operatorname{\mathbb{E}}[\psi^{\prime}(c_{1})]^{2}\sum_{i,j=1}^{n}\mathbf{A}_{n}^{2}(i,j)\leq n\alpha_{n}.$
$\operatorname{\mathbb{E}}R_{3n}=\sum_{i=1}^{n}\operatorname{\mathbb{E}}\left(\sum_{j=1}^{n}\mathbf{A}_{n}(i,j)\psi^{\prime}(c_{j})\right)^{4}\lesssim\sum_{i=1}^{n}\left(\sum_{j=1}^{n}\mathbf{A}_{n}^{2}(i,j)\right)^{2}\leq n\alpha_{n}^{2}.$
@@
We will show that the second moment of the both terms converge to 00. For the first term, we have
$\;\;\;\;\operatorname{\mathbb{E}}\Big{[}\sum_{\mathnormal{i,j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{i}(\psi^{\prime\prime}(c_{i})-\upsilon_{n})\psi^{\prime}(c_{j})\Big{]}^{2}$
$=\sum_{i,j,k,l}\mathbf{A}_{n}(i,j)\mathbf{A}_{n}(k,l)q_{i}q_{k}\operatorname{\mathbb{E}}\left[(\psi^{\prime\prime}(c_{i})-\upsilon)(\psi^{\prime\prime}(c_{k})-\upsilon)\psi^{\prime}(c_{j})\psi^{\prime}(c_{l})\right]$
$\lesssim\sum_{i\neq j}\mathbf{A}_{n}(i,j)^{2}\bigg{[}q_{i}^{2}\operatorname{\mathbb{E}}\left[(\psi^{\prime\prime}(c_{i})-\upsilon)^{2}\psi^{\prime}(c_{j})^{2}\right]+|q_{i}q_{j}|\operatorname{\mathbb{E}}\left[\psi^{\prime}(c_{i})(\psi^{\prime\prime}(c_{i})-\upsilon)\psi^{\prime}(c_{j})(\psi^{\prime\prime}(c_{j})-\upsilon)\right]\bigg{]}$
@@
$\operatorname{\mathbb{E}}\bigg{[}(\upsilon-\upsilon_{n})\sum_{i,j=1}^{n}\mathbf{A}_{n}(i,j)q_{i}\psi^{\prime}(c_{j})\bigg{]}^{2}$ $\leq{\operatorname{\mathbb{E}}(\upsilon-\upsilon_{n})^{2}}{\operatorname{\mathbb{E}}\left(\sum_{i,j=1}^{n}\mathbf{A}_{n}(i,j)q_{i}\psi^{\prime}(c_{j})\right)^{2}}$
@@
$\operatorname{\mathbb{E}}\big{|}\bm{\epsilon}^{\top}\Psi^{\prime}(\mathbf{c})\big{|}\leq\sqrt{\operatorname{\mathbb{E}}\left(\sum_{i=1}^{n}\epsilon_{i}\psi^{\prime}(c_{i})\right)^{2}}\leq\lVert\bm{\epsilon}\rVert.$
@@
$np_{n}\alpha_{n}=\max_{i\in[n]}\frac{d_{i}}{np_{n}}\stackrel{{\scriptstyle P}}{{\to}}1,$
and so $\alpha_{n}=O_{P}\Big{(}\frac{1}{{np_{n}}}\Big{)}\ll n^{-1/2}$.
@@
    With $\mathbf{q}=n^{-1/2}\mathbf{1}$ and $\lambda=1$, we have $(\mathbf{A}_{n}\mathbf{q}-\lambda\mathbf{q})_{i}=\frac{\theta}{\sqrt{n}}\Big{(}\frac{d_{i}}{np_{n}}-1\Big{)}$, we have
@@
It is immediate to check that $\alpha_{n}=d_{n}^{-1}$ and hence $\sqrt{n}\alpha_{n}\leq d_{n}^{-1}\sqrt{n}\to 0$. Also we have $\|{\bf A}_{n}\|_{\infty}=|\theta|$, and so the SHT condition holds for $\theta\in(-1,1)$.
@@
    In this case follows from 2.6 by noting that $(\mathbf{q},\lambda_{n})=(n^{-1/2}\mathbf{1},1)$ forms an exact eigenvector-eigenvalue pair.

### v2

@@ 3 Proof of main results
    where the hidden constants depend on $\kappa,\rho$. Part (b) now follows on plugging-in the bounds in (i) (2.13) in 2.4, and (ii) part (a) of this theorem.
@@
$\operatorname{\mathbb{E}}_{F}(\upsilon_{n}-\upsilon)^{2}$ $=O(\|\mathbf{q}\|_{\infty}^{2}),\quad\operatorname{\mathbb{E}}_{F}R_{1n}=O(\alpha_{n}+\|{\bm{\epsilon}}\|^{2}),\quad\operatorname{\mathbb{E}}_{F}R_{2n}=O(n\alpha_{n}),$ (4.1)
$\quad\operatorname{\mathbb{E}}_{F}R_{3n}$ $=O(n\alpha_{n}^{2}),\quad\operatorname{\mathbb{E}}_{F}R_{4n}=O(\sqrt{\operatorname{\alpha_{\mathnormal{n}}}}+\|\mathbf{q}\|_{\infty}),\quad\operatorname{\mathbb{E}}_{F}|{\bm{\epsilon}}^{\top}\Psi^{\prime}({\bf c})|=O(\|\bm{\epsilon}\|).$
We defer the proof of (4.1) to the end of the proof. In (4.1) and throughout this proof, the implicit constants in $O(\cdot)$ and $\lesssim$ may depend on $\rho$ from (2.2) and $\upsilon$.
@@
    Using 2.4, on the set $\upsilon_{n}\geq\kappa:=\upsilon/2$, the first term in the RHS of (4.2) can be bounded as follows:
    $d_{KS}(T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}),W_{n}\mid{\bf c})\lesssim\sqrt{R_{1n}}+\sqrt{\alpha_{n}R_{2n}}+\sqrt{R_{3n}}+\sqrt{n}\alpha_{n}+\|{\bf q}\|_{\infty}+\|{\bm{\epsilon}}\|=:S^{(1)}_{n}$. (4.3)
@@
    where we use Lemma 3.4, along with the fact that $1-\lambda\upsilon_{n}\in[1-\rho,1+\rho]$. The desired conclusion follows by combining the above three displays, on using (4.1) to note that
    $\operatorname{\mathbb{E}}_{F}[S^{(1)}_{n}]\lesssim\textsc{Err}(\mathbf{A}_{n},\mathbf{q}),\quad\text{ and }\quad\operatorname{\mathbb{P}}(\upsilon_{n}\geq\kappa)\to 1.$
@@
    $X=\mathbf{q}^{\top}\left(\operatorname{\mathnormal{\bm{\sigma}}}-\frac{\Psi^{\prime}(\mathbf{c})}{1-\lambda\upsilon_{n}}\right),~\tau=\frac{\upsilon}{1-\lambda\upsilon},~a=1,~b=\mathbf{q}^{\top}\left(\frac{\Psi^{\prime}(\mathbf{c})}{1-\lambda\upsilon_{n}}-\frac{\Psi^{\prime}(\mathbf{c})}{1-\lambda\upsilon}\right)$
@@
    where in the last inequality we use 2.5 part (b), and restrict ourselves to the set $\upsilon_{n}\geq\kappa=\upsilon/2$, and
    $S^{(2)}_{n}:=\sqrt{R_{2n}}\left(\sqrt{R_{1n}}+\sqrt{\alpha_{n}}+\lVert\bm{\epsilon}\rVert\right)+\sqrt{R_{1n}}+\sqrt{R_{3n}}+\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}+\lVert\mathbf{q}\rVert_{\infty}+R_{4n}+\big|\bm{\epsilon}^{\top}\Psi^{\prime}(\mathbf{c})\big|+\|\bm{\epsilon}\|.$
@@
    $\operatorname{\mathbb{E}}_{F}[S^{(2)}_{n}]\lesssim\textsc{Err}(\mathbf{A}_{n},\mathbf{q})+\sqrt{n\alpha_{n}}\|{\bm{\epsilon}}\|.$
@@
    $d_{KS}\left(\sum_{i=1}^{n}q_{i}\Big(\sigma_{i}-\frac{\psi^{\prime}(c_{i})}{1-\lambda\upsilon}\Big),W_{\infty}\mid{\bf c}\right)=O_{P}(\textsc{Err}(\mathbf{A}_{n},\mathbf{q})+\sqrt{n\alpha_{n}}\|{\bm{\epsilon}}\|),$
@@
    Proceeding to verify the second (i.e. annealed) conclusion of part (b), let (W,W~)($_{\infty},\widetilde{W}_{\infty}$) be mutually independent Gaussians, coupled in the same space as $(\operatorname{\mathnormal{\bm{\sigma}}},{\bf c})$ via an independent coupling. Then, setting $W^{*}_{n}:=\frac{1}{1-\lambda\upsilon}\sum_{i=1}^{n}q_{i}\psi^{\prime}(c_{i})$, triangle inequality gives
[…]
@@
    where the last inequality holds on the set $\upsilon_{n}\geq\kappa=\upsilon/2$, and we have used ((b)). Noting that $d_{KS}$ is always bounded by $1$, the above two displays give
[…]
@@
    $\operatorname{\mathbb{E}}_{F}S_{n}^{(2)}\lesssim$ $\textsc{Err}(\mathbf{A}_{n},\mathbf{q})+\sqrt{n\alpha_{n}}\|{\bm{\epsilon}}\|,$
    $\operatorname{\mathbb{E}}_{F}\left|(\upsilon_{n}-\upsilon)\sum_{i=1}^{n}q_{i}\psi^{\prime}(c_{i})\right|\leq$ $\sqrt{\operatorname{\mathbb{E}}_{F}(\upsilon_{n}-\upsilon)^{2}}\sqrt{\operatorname{\mathbb{E}}_{F}\Big[\sum_{i=1}^{n}q_{i}\psi^{\prime}(c_{i})\Big]^{2}}\lesssim\|{\bf q}\|_{\infty}.$
@@
    $\operatorname{\mathbb{P}}(\upsilon_{n}<\kappa)\leq\operatorname{\mathbb{P}}(|\upsilon_{n}-\upsilon|>\upsilon/2)\lesssim\operatorname{\mathbb{E}}_{F}|\upsilon_{n}-\upsilon|\lesssim\|{\bf q}\|_{\infty},$
    where the last bound again uses (4.1). Finally, the fourth term in the RHS of (4.6) can be bounded using the standard Berry-Esseen CLT for i.i.d. random variables, which gives
@@
Proof of (4.1). First we bound $\upsilon_{n}-\upsilon$. We use the independence of $\mathbf{c}=(c_{1},\ldots,c_{n})^{\top}$ and the definition of $\upsilon=\operatorname{\mathbb{E}}_{F}[\psi^{\prime\prime}(c_{1})]>0$ to get
$\operatorname{\mathbb{E}}_{F}(\upsilon-\upsilon_{n})^{2}=\mbox{Var}(\psi^{\prime\prime}(c_{1}))\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{4}\leq\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{4}\leq\|\mathbf{q}\|_{\infty}^{2}.$ (4.7)
@@
$\operatorname{\mathbb{E}}_{F}R_{1n}$ $=\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\operatorname{\mathbb{E}}_{F}\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}(\psi^{\prime\prime}(c_{j})-\upsilon_{n})\right)^{2}$
$\leq 2\operatorname{\mathbb{E}}_{F}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}(\psi^{\prime\prime}(c_{j})-\upsilon)\right)^{2}+2\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}\right)^{2}\operatorname{\mathbb{E}}_{F}(\upsilon-\upsilon_{n})^{2}$
$=2\operatorname{Var}(\psi^{\prime\prime}(c_{1}))\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)^{2}q_{j}^{2}+2\lVert\mathbf{A}_{n}\mathbf{q}\rVert^{2}\operatorname{\mathbb{E}}_{F}(\upsilon-\upsilon_{n})^{2}$
@@
$\lVert\mathbf{A}_{n}\mathbf{q}\rVert^{2}\operatorname{\mathbb{E}}_{F}(\upsilon-\upsilon_{n})^{2}$ $\leq 2\lVert\bm{\epsilon}\rVert^{2}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{4}+2\sum_{\mathnormal{i}=1}^{\mathnormal{n}}(\lambda q_{i})^{2}q_{i}^{2}$
@@
The bounds for $R_{2n}$ and $R_{3n}$ also follow similarly by explicit computations leveraging the independence of $\mathbf{c}$ and $\operatorname{\mathbb{E}}_{F}[\psi^{\prime}(c_{1})]=0$, as follows:
$\operatorname{\mathbb{E}}_{F}R_{2n}=\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{F}\left(\sum_{j=1}^{n}\mathbf{A}_{n}(i,j)\psi^{\prime}(c_{j})\right)^{2}=\operatorname{\mathbb{E}}_{F}[\psi^{\prime}(c_{1})]^{2}\sum_{i,j=1}^{n}\mathbf{A}_{n}^{2}(i,j)\leq n\alpha_{n}.$
$\operatorname{\mathbb{E}}_{F}R_{3n}=\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{F}\left(\sum_{j=1}^{n}\mathbf{A}_{n}(i,j)\psi^{\prime}(c_{j})\right)^{4}\lesssim\sum_{i=1}^{n}\left(\sum_{j=1}^{n}\mathbf{A}_{n}^{2}(i,j)\right)^{2}\leq n\alpha_{n}^{2}.$
@@
We will show that the second moment of the both terms converge to 0. For the first term, we have
$\;\;\;\;\operatorname{\mathbb{E}}_{F}\Big[\sum_{\mathnormal{i,j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{i}(\psi^{\prime\prime}(c_{i})-\upsilon)\psi^{\prime}(c_{j})\Big]^{2}$
$=\sum_{i,j,k,l}\mathbf{A}_{n}(i,j)\mathbf{A}_{n}(k,l)q_{i}q_{k}\operatorname{\mathbb{E}}_{F}\left[(\psi^{\prime\prime}(c_{i})-\upsilon)(\psi^{\prime\prime}(c_{k})-\upsilon)\psi^{\prime}(c_{j})\psi^{\prime}(c_{l})\right]$
$\lesssim\sum_{i\neq j}\mathbf{A}_{n}(i,j)^{2}\bigg[q_{i}^{2}\operatorname{\mathbb{E}}_{F}\left[(\psi^{\prime\prime}(c_{i})-\upsilon)^{2}\psi^{\prime}(c_{j})^{2}\right]+|q_{i}q_{j}|\operatorname{\mathbb{E}}_{F}\Big|\psi^{\prime}(c_{i})(\psi^{\prime\prime}(c_{i})-\upsilon)\psi^{\prime}(c_{j})(\psi^{\prime\prime}(c_{j})-\upsilon)\Big|\bigg]$
@@
$\operatorname{\mathbb{E}}_{F}\bigg[(\upsilon-\upsilon_{n})\sum_{i,j=1}^{n}\mathbf{A}_{n}(i,j)q_{i}\psi^{\prime}(c_{j})\bigg]^{2}$ $\leq{\operatorname{\mathbb{E}}_{F}(\upsilon-\upsilon_{n})^{2}}{\operatorname{\mathbb{E}}_{F}\left(\sum_{i,j=1}^{n}\mathbf{A}_{n}(i,j)q_{i}\psi^{\prime}(c_{j})\right)^{2}}$
@@
$\operatorname{\mathbb{E}}_{F}\big|\bm{\epsilon}^{\top}\Psi^{\prime}(\mathbf{c})\big|\leq\sqrt{\operatorname{\mathbb{E}}_{F}\left(\sum_{i=1}^{n}\epsilon_{i}\psi^{\prime}(c_{i})\right)^{2}}\leq\lVert\bm{\epsilon}\rVert.$
@@
$np_{n}\alpha_{n}=\theta^{2}\max_{i\in[n]}\frac{d_{i}}{np_{n}}\stackrel{{\scriptstyle P}}{{\to}}\theta^{2},$
and so $\alpha_{n}=O_{P}\Big(\frac{1}{{np_{n}}}\Big)=o_{P}(n^{-1/2})$.
@@
    With $\mathbf{q}=n^{-1/2}\mathbf{1}$ and $\lambda=\theta$, we have $(\mathbf{A}_{n}\mathbf{q}-\lambda\mathbf{q})_{i}=\frac{\theta}{\sqrt{n}}\Big(\frac{d_{i}}{np_{n}}-1\Big)$, we have
@@
For simplicity, we only show the convergence of two dimensional marginals for part (a) of 2.8. The argument directly generalizes to part (b), and we omit the details.
Write $\upsilon:=\operatorname{\mathbb{E}}\psi^{\prime\prime}(c_{1})$. By the Cramér–Wold device, it suffices to show that for any $a,b\in\mathbb{R}$ and $Q_{1},Q_{2}\in\mathcal{Q}$,
$a\sqrt{n}\langle Q_{1},g_{\operatorname{\mathnormal{\bm{\sigma}}}}-g_{{\bf u}}\rangle+b\sqrt{n}\langle Q_{2},g_{\operatorname{\mathnormal{\bm{\sigma}}}}-g_{{\bf u}}\rangle\mid\mathbf{c}\operatorname*{\xrightarrow{\mathnormal{d}}}\operatorname{N}\Big(0,\upsilon\underbrace{(a^{2}+2ab\int_{0}^{1}Q_{1}(x)Q_{2}(x)+b^{2})}_{:=F_{a,b}(Q_{1},Q_{2})}\Big).$
Setting $Q(x):=\frac{aQ_{1}(x)+bQ_{2}(x)}{\sqrt{F_{a,b}(Q_{1},Q_{2})}}$ we then have $Q\in\mathcal{Q}$, and
$\sqrt{n}\langle Q,g_{\operatorname{\mathnormal{\bm{\sigma}}}}-g_{{\bf u}}\rangle=a\frac{\sqrt{n}\langle Q_{1},g_{\operatorname{\mathnormal{\bm{\sigma}}}}-g_{{\bf u}}\rangle}{\sqrt{F_{a,b}(Q_{1},Q_{2})}}+b\frac{\sqrt{n}\langle Q_{2},g_{\operatorname{\mathnormal{\bm{\sigma}}}}-g_{{\bf u}}\rangle}{\sqrt{F_{a,b}(Q_{1},Q_{2})}}.$
It thus suffices to show that for all $Q\in\mathcal{Q}$ we have
$\sqrt{n}\langle Q,g_{\operatorname{\mathnormal{\bm{\sigma}}}}-g_{{\bf u}}\rangle\mid{\bf c}\stackrel{{\scriptstyle d}}{{\to}}N(0,\upsilon).$
Setting
$q_{i}:=\sqrt{n}\int_{I_{i}}Q(x)dx\qquad\text{ where }\qquad I_{i}:=\Big(\frac{i-1}{n},\frac{i}{n}\Big],$
the above is equivalent to the claim
$\sum_{i=1}^{n}q_{i}(\sigma_{i}-u_{i})\mid{\bf c}\stackrel{{\scriptstyle d}}{{\to}}N(0,\upsilon).$
This follows by invoking part (b) of 2.7, for which we need to verify that $\bf q$} satisfies (2.19), part (b) of 2.7 implies the desired convergence. We check each parts of (2.19) below.
  • $\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}\to 1$
    Let $U$ be a random variable which is uniform on $(0,1]$, and let $\mathcal{F}_{n}$ be a $\sigma$-field defined by $\mathcal{F}_{n}:=\sigma(\{U\in I_{i}:i\in[n]\})$. Then, setting
    $Y_{n}:=\operatorname{\mathbb{E}}[Q(U)\mid\mathcal{F}_{n}]=1\{U\in I_{i}\}n\int_{I_{i}}Q(x)dx=1\{U\in I_{i}\}\sqrt{n}q_{i},$
    by Lévy’s zero-one law (e.g. Theorem 4.6.8 in [26]) we get $Y_{n}\stackrel{{\scriptstyle L^{2}}}{{\to}}Q(U)$, and consequently
    $\sum_{i=1}^{n}q_{i}^{2}=\operatorname{\mathbb{E}}Y_{n}^{2}\to\operatorname{\mathbb{E}}Q(U)^{2}=\int_{0}^{1}Q(x)^{2}dx=1,$
    as desired. Note that while (2.19) requires the equality $\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}=1$, this can be achieved by re-normalizing $\mathbf{q}$ without changing the limiting variance.
  • $\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}=0$
    This follows on noting that $\sum_{i=1}^{n}q_{i}=\sqrt{n}\int_{0}^{1}Q(x)dx=0$.
  • $\|\mathbf{q}\|_{\infty}\to 0$
    This follows on noting that
    $\|\mathbf{q}\|_{\infty}=\max_{\mathnormal{i}=1}^{\mathnormal{n}}\sqrt{n}\int_{I_{i}}Q(x)dx\leq\max_{\mathnormal{i}=1}^{\mathnormal{n}}\sqrt{\int_{I_{i}}Q(x)^{2}dx}\to 0,$
    where the last limit follows on using DCT.
Proof of 2.9.
It is immediate to check that $\alpha_{n}=\theta^{2}/d(n)$ and hence $\sqrt{n}\alpha_{n}=\theta^{2}\sqrt{n}/d(n)\to 0$. Also we have $\|{\bf A}_{n}\|_{\infty}=|\theta|$, and so the SHT condition holds for $\theta\in(-1,1)$.
@@
    In this case follows from 2.6 by noting that $(\mathbf{q},\lambda_{n})=(n^{-1/2}\mathbf{1},\theta)$ forms an exact eigenvector-eigenvalue pair.